# Matrix-Driven Instant Review: Confident Detection and Reconstruction of LLM Plagiarism on PC

### v2

@@ 3.3 Equivalent Transformations
where $\otimes$ denotes the tensor product of matrices, $\mu\neq 0$ is a scalar, $P_{1}\in\mathrm{Perm}(\mathrm{NumHeads})$ is a permutation matrix over $\mathrm{NumHeads}$ channels, $P_{2}\in\mathrm{Perm}(\mathrm{QueriesPerHead})$ is a permutation matrix over $\mathrm{QueriesPerHead}$ queries, $S\in\mathrm{diag}(\pm 1,\cdots,\pm 1)\in\mathrm{M}_{\mathrm{HeadDim}}(\mathbb{R})$, and $H_{v}\in\mathrm{GL}(\mathrm{HeadDim},\mathbb{R})$ is an arbitrary invertible matrix.
This set does not account for all possible configurations. For example, $S$ could be replaced by an arbitrary orthogonal matrix. However, such substitutions might obfuscate the implementation of Rotary Positional Embeddings (RoPE) [43], potentially flagging the model for plagiarism. To ensure compatibility with both QK-norm [20] and RoPE, we restrict $S$ to diagonal matrices with entries $\pm 1$:
$S\in\mathrm{diag}(\pm 1,\cdots,\pm 1).$
In practice, these transformation matrices may all reduce to identity matrices, particularly when the victim and suspect models are just identical.
3.3.1 Transformations with Training and Noise Injection
After applying all equivalent transformations, Model $B$ still inherits its policy entirely from Model $A$, i.e., $\pi(B)=\pi(A)$. As a result, $B$ remains easily detectable via retrieval-based watermarks, KL divergence, or simple output comparisons.
To obscure these patterns, the model stealer might train the model on a large number of tokens. In such cases, our formulas can be adjusted as follows:
$Q^{\prime}$ $=U_{Q}QW_{Q}+N_{Q},$
$K^{\prime}$ $=U_{K}KW_{K}+N_{K},$
$V^{\prime}$ $=U_{V}VW_{V}+N_{V},$
$O^{\prime}$ $=W_{O}^{-1}OU_{O}^{-1}+N_{O},$
where $N_{Q},N_{K},N_{V},N_{O}$ are noise matrices. We assume that the noise has a very small root-mean-square (RMS) magnitude compared to the original matrices $Q,K,V,O$.
3.3.2 Pruning and Upcycling
Without loss of generality, we may assume that Model $B$ is a pruned version of Model $A$. This formulation is analogous to Model $A$ being upcycled from Model $B$.
For simplicity, we provide a possible relationship:
$Q^{\prime}$ $=U_{Q}QW_{Q}+N_{Q},$
$K^{\prime}$ $=U_{K}KW_{K}+N_{K},$
$V^{\prime}$ $=U_{V}VW_{V}+N_{V},$
$O^{\prime}$ $=W_{O}OU_{O}^{T}+N_{O},$
where $U_{Q}=U_{K}=U_{V}=U_{O}=U$ is the same partial isometry (i.e., $UU^{T}=\bm{1}$); $W_{Q}$ and $W_{K}$ are also partial isometries satisfying $W_{Q}^{T}W_{Q}=\bm{1}$ and $W_{K}^{T}W_{K}=\bm{1}$. The possible forms of $W_{V}$ and $W_{O}$ involving pruning can be highly complex. We leave their exact characterization for future work.
3.4 Preliminary Examinations
To effectively identify model plagiarism, we must provide sufficient evidence and even reconstruct the entire fabrication process. To achieve this goal, we must systematically solve all permutations and transformations involved.
Previous research [53] correctly identified the following covariant relationships:
$Q^{\prime}K^{\prime T}$ $\approx UQK^{T}U^{T},$
$V^{\prime}O^{\prime}$ $\approx UVOU^{T},$
where $U$ is an orthogonal transformation matrix. Since there is only one matrix $U$ to be solved, it is feasible to estimate $U$ by minimizing the error:
$\min_{U\in\mathrm{O}(\mathrm{EmbDim})}\left(\lVert Q^{\prime}K^{\prime T}-UQK^{T}U^{T}\rVert_{F}^{2}+\lVert V^{\prime}O^{\prime}-UVOU^{T}\rVert_{F}^{2}\right).$
Because $U$ represents the same transformation across all layers, we can extend this optimization to all layers:
$\min_{U\in\mathrm{O}(\mathrm{EmbDim})}\sum_{l=1}^{L}\left(\lVert Q^{\prime}_{l}K^{\prime T}_{l}-UQ_{l}K^{T}_{l}U^{T}\rVert_{F}^{2}+\lVert V^{\prime}_{l}O^{\prime}_{l}-UV_{l}O_{l}U^{T}\rVert_{F}^{2}\right),$
or equivalently,
$\max_{U\in\mathrm{O}(\mathrm{EmbDim})}\sum_{l=1}^{L}\left(2\mathrm{Tr}(UQ_{l}K^{T}_{l}U^{T}K^{\prime}_{l}Q^{\prime T}_{l})+2\mathrm{Tr}(UV_{l}O_{l}U^{T}O^{\prime T}_{l}V^{\prime T}_{l})\right).$
Despite systematic approaches for maximizing a single term of $\mathrm{Tr}(UAU^{T}B)$ [8], this optimization problem involving multiple such terms is challenging and may not admit a unique solution. For example: If $U$ is a solution, then U-U is also a solution. In extreme cases, if all $Q_{l}K^{T}_{l}$, $Q^{\prime}_{l}K^{\prime T}_{l}$, $V^{\prime}_{l}O^{\prime}_{l}$, and $V_{l}O_{l}$ are identity matrices, then any $U\in\mathrm{O}(\mathrm{EmbDim})$ satisfies the equations.
3.5 Breaking the Shell: Vocabulary Embeddings
Let $E\in\mathbb{R}^{\mathrm{VocabSize}\times\mathrm{EmbDim}}$ denote the vocabulary embedding matrix of model $A$, and $E^{\prime}\in\mathbb{R}^{\mathrm{VocabSize}\times\mathrm{EmbDim}}$ for model $B$. Following the conventions established in previous discussions, we have:
$E^{\prime}=EU^{T}+N_{E},$
To minimize the difference between $E^{\prime}$ and $EX^{T}$, we solve the following optimization problem:
$\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\lVert E^{\prime}-EX^{T}\rVert_{F}^{2}=\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\left\langle E^{\prime}-EX^{T},E^{\prime}-EX^{T}\right\rangle_{F}.$
Expanding the Frobenius norm yields:
$\arg\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\lVert E^{\prime}-EX^{T}\rVert_{F}^{2}$ $=\arg\min_{X\in\mathrm{O}(\mathrm{EmbDim})}\left(\lVert E^{\prime}\rVert_{F}^{2}+\lVert E\rVert_{F}^{2}-2\left\langle EX^{T},E^{\prime}\right\rangle_{F}\right)$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\left\langle EX^{T},E^{\prime}\right\rangle_{F}$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left(EX^{T}E^{\prime T}\right)$
$=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left((E^{\prime T}E)X^{T}\right).$
We denote $\tilde{U}$ as the solution to this optimization problem, which corresponds to the orthogonal part of $(E^{\prime T}E)$. Note that $\tilde{U}$ is not the ground truth of $U$, but rather a close approximation. If $\tilde{U}$ is sufficiently close to a permutation matrix, we can identify a permutation matrix $P\in\mathrm{Perm}(\mathrm{EmbDim})$ via:
$P=\arg\max\operatorname{Tr}(P\tilde{U}^{T}).$
This problem is equivalent to solving maximum bipartite matching or the linear sum assignment problem [9], which can be computed in $O(n^{3})$ time using the Hungarian algorithm. For large matrices (e.g., $18432\times 18432$), solving this exactly may take approximately one hour on a CPU. To reduce computational cost, we propose a faster heuristic: compute the positions of the maximal element in each row and verify whether they form a valid permutation.
A unique determination of $\tilde{U}$ requires $(E^{\prime T}E)$ to be non-degenerate and full-rank, which necessitates the vocabulary size to satisfy $\mathrm{VocabSize}\geq\mathrm{EmbDim}$. This condition is easily satisfied, as most recent tokenizers have $\mathrm{VocabSize}\geq 3\times 10^{4}$.
3.5.1 When the Tokenizer Is Changed
If the model stealer attempts to obfuscate the embeddings by using a different tokenizer, the situation becomes more complex. However, we expect a significant overlap in tokens between model $A$’s and model $B$’s tokenizer, including:
  • Single-byte ASCII characters (the smallest unit in byte-pair encoding [38]);
  • Common English words such as am, is and take;
  • Accented letters like é.
These tokens collectively constitute a large subset of the vocabulary, and we anticipate the size of their intersection to exceed $1\times 10^{4}$.
Drawing from Saussure’s theory of linguistics [36], tokens are arbitrary, but their contextual usage systematically reveals their meaning. Based on this assumption, even after reinitialization, the embeddings of shared tokens, trained on billions of words, should gradually align.
Let $\mathcal{C}$ denote the set of all common tokens. We estimate $\tilde{U}$ as:
$\tilde{U}=\arg\max_{X\in\mathrm{O}(\mathrm{EmbDim})}\operatorname{Tr}\left(\left(E^{\prime}[\mathcal{C},:]^{T}E[\mathcal{C},:]\right)X^{T}\right).$
Thus, $\tilde{U}$ corresponds to the orthogonal part in the polar decomposition of $(E^{\prime}[\mathcal{C},:]^{T}E[\mathcal{C},:])$.
3.6 Estimating the $p$-Value
We aim to identify a permutation matrix $P$ from $\tilde{U}$ as $P=\arg\max\operatorname{Tr}(P\tilde{U}^{T})$. However, we need a statistical criterion to determine whether our identification is successful. While it may be reasonable to visually inspect plots of $P$ and $\tilde{U}$, the value of $\operatorname{Tr}(P\tilde{U}^{T})$ itself often provides strong evidence. In some experiments, this value can reach as high as $0.2\cdot\mathrm{EmbDim}$ to $0.99\cdot\mathrm{EmbDim}$.
Null Hypothesis. Our null hypothesis assumes that models $A$ and $B$ are not homologous, and there is no apparent similarity between their weights. Under this assumption, the probability measure $\mathrm{d}\mathbb{P}$ should be uniform across all admissible transformations. For instance, for the vocabulary embedding matrix $E$ of model $A$, its density satisfies:
$E\mathrm{d}\mathbb{P}=(EG)\mathrm{d}\mathbb{P},\quad\forall G\in\mathrm{O}(n),$
where $G$ is an orthogonal matrix. Consequently, the distribution of $\tilde{U}$, as the orthogonal part of $(E^{\prime T}E)$, satisfies:
$\tilde{U}\mathrm{d}\mathbb{P}=(\tilde{U}G)\mathrm{d}\mathbb{P}.$
This implies that $\tilde{U}$ is uniformly distributed over $\mathrm{O}(n)$ according to the Haar measure.
Now, fix $P_{0}$ as an arbitrary permutation matrix. We estimate the probability of $\operatorname{Tr}(P_{0}\tilde{U}^{T})\geq c$, denoted as:
$f(c):=\mathbb{P}\left[\operatorname{Tr}(P_{0}\tilde{U}^{T})\geq c\right].$
Since $\tilde{U}^{T}$ is uniformly distributed under the Haar measure, we have:
$(P_{0}\tilde{U}^{T})\mathrm{d}\mathbb{P}=(P_{0}P_{0}^{T}\tilde{U}^{T})\mathrm{d}\mathbb{P}=\tilde{U}^{T}\mathrm{d}\mathbb{P}.$
Thus:
@@ 3.6 Estimating the $p$-Value
However, our permutation $P$ maximizes $\operatorname{Tr}(P\tilde{U}^{T})$. With $n!$ possible permutations, only the one maximizing $\operatorname{Tr}(P\tilde{U}^{T})$ is chosen. This implies that the $p$-value satisfies:
$p$ $=\max_{P\in\mathrm{Perm}(n)}\mathbb{P}\left[\operatorname{Tr}(P\tilde{U})\geq c\right]$
$=\mathbb{P}\left[\exists P\in\mathrm{Perm}(n),\ \operatorname{Tr}(P\tilde{U})\geq c\right]$
$\leq\sum_{P_{0}\in\mathrm{Perm}(n)}\mathbb{P}\left[\operatorname{Tr}(P_{0}\tilde{U})\geq c\right]$
$=n!\cdot f(c).$
This result indicates that we can estimate the $p$-value based on the evaluation of $f(c)$.
@@
To study the tail behavior of $f(c)$, we leverage tools from Large Deviation Theory. Since $n$ is typically large and the embedding dimension $n$ is usually an even number in most models, we assume $n=2m$ without loss of generality. Additionally, we assume $\det(\tilde{U})=1$, or equivalently $\tilde{U}\in\mathrm{SO}(2m)$, since this introduces only an infinitesimal difference in the thermodynamic limit ($n\to\infty$; see Appendix A). These assumptions simplify our analysis while preserving accuracy.

### v3

@@ Appendix E Inner Transformations in the Attention Module
where $\otimes$ denotes the Kronecker product of matrices, $\mu\neq 0$ is a scalar, $P_{1}\in\mathrm{Perm}(\mathrm{NumKVHeads})$ is a permutation matrix over $\mathrm{NumKVHeads}$ channels (and $\sigma$ is the corresponding permutation such that $\sum\mathbf{1}_{v,\sigma(v)}=P_{1}$), $P_{2}\in\mathrm{Perm}(\mathrm{QueriesPerHead})$ is a permutation matrix over $\mathrm{QueriesPerHead}$ queries, $S\in\mathrm{diag}(\pm 1,\cdots,\pm 1)\in\mathrm{M}_{\mathrm{HeadDim}}(\mathbb{R})$, and $H\in\mathrm{O}(\mathrm{HeadDim},\mathbb{R})$ is an arbitrary orthogonal matrix.
To ensure compatibility with both QK-norm (henry2020querykeynormalizationtransformers) and RoPE (su2023roformerenhancedtransformerrotary), $S$ is restricted to diagonal matrices with entries $\pm 1$: $S\in\mathrm{diag}(\pm 1,\cdots,\pm 1).$
The totally invariant group $G$ can be generated by these elements:
$U,\ \ (P_{1,\ell})_{1\leq\ell\leq L},\ \ (P_{2,\ell})_{1\leq\ell\leq L},\ \ (S_{\ell})_{1\leq\ell\leq L},$
$(H_{\ell,v})_{1\leq\ell\leq L,1\leq v\leq\mathrm{NumKVHeads}}.$
Where $U\in\mathrm{O}(\mathrm{EmbDim})$ is the component that contributes the most dimension ($(\mathrm{EmbDim}^{2}-\mathrm{EmbDim})/2$) to this group.
Each component satisfies the constraints outlined in the paper:
  • $U$ represents the outer transformation, which is shared across all layers and belongs to the orthogonal group $\mathrm{O}(\mathrm{EmbDim})$.
  • $(P_{1,\ell})_{1\leq\ell\leq L}$ and $(P_{2,\ell})_{1\leq\ell\leq L}$ are permutation matrices corresponding to attention heads and queries respectively.
  • $(S_{\ell})_{1\leq\ell\leq L}$ are diagonal matrices with entries $\pm 1$ to ensure compatibility with Rotary Positional Embeddings.
  • $(H_{\ell,v})_{1\leq\ell\leq L,1\leq v\leq\mathrm{NumKVHeads}}$ are general invertible matrices applied to individual attention heads.
In other words, $G$ is the Cartesian product of these components:
$G=(U)\times\prod_{\ell=1}^{L}\left((P_{1,\ell})\times(P_{2,\ell})\times(S_{\ell})\times\prod_{v=1}^{\mathrm{NumKVHeads}}(H_{\ell,v})\right)$
We summarize the GQA architecture and invariant transformations in Figure 6, illustrated using Penrose tensor notation, which compactly encodes the tensor contractions and symmetry operations underlying the GQA module, and may serve a visual illustration why the inner transformations take the form above.
Figure 6: A Penrose notation of a model with Grouped Query Attention architecture (model AA, up) and an illustration of an adversary $B$ under application of invariant transformations (down). Bias terms are omitted, and RMSNorms are not explicitly shown.
Appendix F Solving All Transformations
If the significance of the $p$-value has already been determined from the $p$-values for vocabulary embeddings, this step is entirely optional. However, if the no significant $p$-value is observed at this stage, we cannot yet rule out the possibility of model homology. This may occur when an adversary used a general orthogonal matrix for obfuscation. Additionally, interested readers may wish to determine the exact relationship between two models, potentially with different architectures.
Since $P=\arg\max_{P}\operatorname{Tr}(P\tilde{U}^{\text{T}})$, we proceed as follows: If $P$ is reliably identified as a permutation matrix ($p<p_{0}$, where $p_{0}$ is some threshold), set $U:=P$; otherwise, set $U:=\tilde{U}$.
Solving the Relationship between Layers. In practical scenarios, the number of layers for two models are not necessarily the same. It is important to first determine the relationship between the layers of both models via solving a maximal ordered matching problem. We select a representative matrix (for example, attention K or V) for each layer $i$ of model $A$ and layer $j$ of model $B$. For each pair $(i,j)$ we do the process below and obtain a similarity measure matrix of shape $L\times L^{\prime}$. Solving the maximal ordered matching for this matrix will give the relationship between all layers. See Section I for a specific example how this works.
Our objectives are:
$\min_{W_{Q}\text{ orthogonal}}\lVert UQW_{Q}-Q^{\prime}\rVert_{F}^{2},\ \ \min_{W_{K}\text{ orthogonal}}\lVert UKW_{K}-K^{\prime}\rVert_{F}^{2},$
$W_{Q}=\lambda\mathrm{Ortho}(Q^{\text{T}}U^{\text{T}}Q^{\prime}),\quad W_{K}=\mu\mathrm{Ortho}(K^{\text{T}}U^{\text{T}}K^{\prime}),$
where the layer-wise similarity measure can be calculated as
$t_{\square}=\texttt{linear\_sum\_assignment}(W_{\square}),\quad\square\in Q,K.$
Also, scaling coefficients $\lambda$ and $\mu$ are computed as:
For $W_{V}$ and $W_{O}$, we may apply the same method to compute $\mathrm{Ortho}(V^{\text{T}}U^{\text{T}}V^{\prime})$ and $\mathrm{Ortho}(OU^{\text{T}}O^{\prime\text{T}})$. However, since $W_{V}$ and $W_{O}$ may involve general invertible transformations, our method does not guarantee recovering the exact transformation. solving the general case without additional assumptions is challenging. We leave this problem for future work.
$U_{X}$ $=\mathrm{Ortho}(X^{\text{T}}U^{\text{T}}X^{\prime}),\quad X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\};$
$P$ $=\operatorname*{arg\,max}_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(P\left(U_{\mathrm{Gate}}+U_{\mathrm{Up}}+U_{\mathrm{Down}}\right)^{\text{T}}\right).$
Typically, we expect the three solutions $\arg\max\operatorname{Tr}\left(PU_{X}^{\text{T}}\right)$, for $X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\}$, to yield the same permutation. However, computing the orthonormal part for intermediate matrices (which often have 14,000–20,000 rows) is computationally expensive. Adding the three terms together would triple the computation cost. If the noise level is tolerable, we may simply select one of them:
$P=\operatorname*{arg\,max}_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(PU_{\mathrm{Up}}^{\text{T}}\right).$
Appendix G Overall Algorithm
We summarize our algorithm in three steps: embedding analysis (Algorithm 1), layer correspondence matching (Algorithm 2), and layer analysis (Algorithm 3). Embedding analysis is the first step for MDIR homology detection. In most cases not involving a sophisticated adversary, this is all that is needed to make a strong detection. For model pairs with different numbers of layers, one can then run MDIR layer correspondence matching. This results in a $p$-value for every pair of layers $(i,j)$ which can be used to determine which layers match. Finally, MDIR layer analysis can be run on each pair of matching layers.
Algorithm 2 MDIR layer correspondence matching
Layer $i\in[L_{A}]$Layer $j\in[L_{B}]$$(X,X^{\prime})\in\{V\}$of $(A_{i},B_{j})$$U_{X}\leftarrow\mathrm{Ortho}(X^{\text{T}}U_{E}^{\text{T}}X^{\prime})$$P_{X}\leftarrow\arg\max_{P\in\mathrm{Perm}(\mathrm{D_{attn}})}\operatorname{Tr}(PU_{X}^{\text{T}})$$p_{X}[i,j]\leftarrow L\cdot\mathrm{PValue}(P_{X},U_{X},\mathrm{D_{attn}})$
[…]
Algorithm 3 MDIR layer analysis
[…]
Appendix H Estimation of $p$-value via Large Deviation Theory
Let $\tilde{U}$ be a random orthogonal matrix, distributed uniformly according to the normalized Haar measure. We aim to estimate the following function, especially its long-tail behavior, for $c>0$:
@@ Appendix H Estimation of $p$-value via Large Deviation Theory
To study the tail behavior of $f(c)$, we leverage tools from Large Deviation Theory. Since $n$ is typically large and the embedding dimension $n$ is usually an even number in most models, we assume $n=2m$ without loss of generality. Additionally, we assume $\det(\tilde{U})=1$, or equivalently $\tilde{U}\in\mathrm{SO}(2m)$, since this introduces only an infinitesimal difference in the thermodynamic limit ($n\to\infty$). These assumptions simplify our analysis while preserving accuracy.